OpenStudy (anonymous):

Solve the system, using substitution. 4x + y = 4 and y = 4 - 4x no solution infinitely many solutions

OpenStudy (anonymous):

4x + y = 4 --(I) y = 4 - 4x --(II) If we rearrange (I), we get y= 4 - 4x which is the same as (II) so there will be infinitely many solutions

OpenStudy (anonymous):

By method of substitution, it is as follows:

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

4x + y = 4 ---(I) y = 4 - 4x ---(II) substitute (II) into (I) 4x + (4 - 4x) = 4 4x +4 - 4x = 4 4x and - 4x cancel out leaving 4 = 4 This is a true statement and hence the given system of equations has infinitely many solutions